OpenStudy (anonymous):

HELP PLEASE Solve for x: 5x2 + 25x = 0 (2 points) x = 0, x = 5 x = 0, x = -5 x = -5, x = 0, x = 5 x = 5, x = 25

OpenStudy (arindameducationusc):

5x^2+25x=0 is this the question?

OpenStudy (anonymous):

yes

OpenStudy (arindameducationusc):

very easy.....

OpenStudy (arindameducationusc):

just take out common x

OpenStudy (arindameducationusc):

then equate x=0 and 5x+25=0

OpenStudy (arindameducationusc):

got?

OpenStudy (anonymous):

okay, o it is X=0, X=5

OpenStudy (arindameducationusc):

no x=-5

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

is it bc we are subtracting 5?

OpenStudy (anonymous):

\[5x^2 + 25x = 0\\5x.x+5.5.x=0\\5x(x+5)=0\\ \implies\ either\ 5x=0\ or\ (x+5)=0\\ \implies\ either\ x=0\ or\ x=-5\\\]

OpenStudy (arindameducationusc):

when we shift 25 to the other side it becomes -25....

OpenStudy (anonymous):

okay

OpenStudy (arindameducationusc):

follow @eninone steps its right....

OpenStudy (arindameducationusc):

anyone you can do....

OpenStudy (anonymous):

but there is only X=5 and X=25

OpenStudy (anonymous):

never mind i know now

OpenStudy (arindameducationusc):

Ok....